OpenStudy (anonymous):

(08.06 MC)The distances (y), in miles, of two cars from their starting points at certain times (x), in hours, are shown by the equations below: Car A y = 55x + 32 Car B y = 42x + 58 After how many hours will the two cars be at the same distance from their starting point and what will that distance be?

OpenStudy (anonymous):

@claringleek

OpenStudy (anonymous):

answer choices?

OpenStudy (anonymous):

2 hours, 142 miles 2 hours, 145 miles 3 hours, 142 miles 3 hours, 145 miles

OpenStudy (anonymous):

@claringleek

OpenStudy (anonymous):

ok so b and d are defiantly out the picture im trying to figure out between a and c

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

im going with A

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

can you answer a few more

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

(08.06 LC)Variable c is 9 more than variable w. Variable c is also 3 less than variable w. Which pair of equations best models the relationship between c and w? c = w − 9 c = 4w + 3 c = w + 9 c = w − 3 w = c + 9 w = c − 3 w = c − 9 w = 4c + 3

OpenStudy (mathstudent55):

Since y represents the distance, and you want the two cars to be at the same distance, set the two equations equal to each other, and solve for x. That will give you the number of hours. Then use either of the original equations with the value of x you found and solve for y. That will give you the distance.

OpenStudy (mathstudent55):

The above was for the first problem.

OpenStudy (mathstudent55):

For the second problem: Variable c is 9 more than variable w. How do you write an expression to mean 9 more than w?

OpenStudy (anonymous):

they keep mentioning c so put in what they said w= 9+c and w= 3-c

OpenStudy (anonymous):

thanks guys @claringleek @mathstudent55

OpenStudy (anonymous):

Meg is 6 years older than Victor. Meg's age is 2 years less than five times Victor's age. The equations below model the relationship between Meg's age (m) and Victor's age (v): m = v + 6 m = 5v − 2 Which is a possible correct method to find Meg's and Victor's ages? Solve m + 6 = 5m − 2 to find the value of m. Write the points where the graphs of the equations intersect the x axis. Solve v + 6 = 5v − 2 to find the value of v. Write the points where the graphs of the equations intersect the y axis.

OpenStudy (mathstudent55):

c is 9 more than w is c = w + 9 This equation has to be in the answer. c is 3 less than w is c = w - 3 This equation also has to b in the answer.

OpenStudy (mathstudent55):

@leah.desrosiers Can you please start a new post for each new question?

OpenStudy (anonymous):

so whats the answer? B

OpenStudy (anonymous):

@claringleek